# ROLCH: Regularized Online Learning for Conditional Heteroskedasticity

### v1

@@ 2.1 Batch GAMLSS
        Regress $\boldsymbol{z}_{k}^{[m,r]}$ against the design matrix ${\mathcal{X}}_{k}$ using the iterative weights $\boldsymbol{W}_{kk}^{[m,r]}$ to obtain the parameter estimates $\boldsymbol{\beta}_{k}^{[m,r+1]}$.
      • iii.
        Calculate the updated $\boldsymbol{\eta}_{k}^{[m,r+1]}$ and $\boldsymbol{\theta}_{k}^{[m,r+1]}$.
      • iv.
        Evaluate the convergence.
    2. (b)
      End the inner cycle on the convergence of $\boldsymbol{\beta}_{k}^{[m,...]}$ and set $\boldsymbol{\beta}_{k}^{[m+1,0]}=\boldsymbol{\beta}_{k}^{[m,...]}$, set $\boldsymbol{\eta}_{k}^{[m+1,0]}=\boldsymbol{\eta}_{k}^{[m,...]}$ and set $\boldsymbol{\theta}^{[m+1,0]}=\boldsymbol{\theta}_{k}^{[m,...]}$. Otherwise, continue the inner cycle until convergence.
  3. 3.
    (Next Parameter:) Update $k$ upon convergence.
  4. 4.
    (End:) End the outer cycle if the change in the penalized likelihood is sufficiently small, otherwise update $m$ and continue the outer cycle.
Note that step 2 – b is the core estimation of the conditional distribution parameters and that the RS algorithm is agnostic to which (weighted) statistical learning method employed to regress $\boldsymbol{z}_{k}^{[m,r]}$ against the design matrix ${\mathcal{X}}_{k}$ using the iterative weights $\boldsymbol{W}_{kk}^{[m,r]}$. This has facilitated the development of multiple extensions to the batch GAMLSS, such as gamlss.boost (Mayr et al., 2012; Hofner et al., 2014) or regularized approaches as in Groll et al. (2019) and gamlss.lasso (Ziel et al., 2021). We will employ this agnosticism and use online coordinate descent (OCD) to update the distributional parameters online.
Relation to iteratively reweighted least squares
As an aside, it is possible to show that the RS algorithm, using the Newton-Rapson scoring, nests the iteratively reweighted least squares algorithm (IRLS) for conditional heteroskedasticity (Ziel et al., 2016; Ziel, 2016) if $\mathcal{D}$ is chosen as the mean-variance parameterization of the normal distribution and both $g_{0}(\cdot)$ and $g_{1}(\cdot)$ are the identity link functions.
Propagation of errors
Note that, as in many instances of distributional modelling, the model for the conditional mean $\hat{\mu}={\mathcal{X}}_{\mu}\hat{\beta}_{\mu}$ is part of the model for the conditional variance due to the relation $\sigma^{2}=\left({\mathcal{Y}}-\hat{\mu}\right)^{2}$. Therefore, a low-quality model for $\hat{\mu}$ will propagate through the estimation and overestimate the conditional variance and potentially higher moments (Ziel, 2022).
@@ 2.2 Online estimation of regularized linear models
Let us now move to the online or streaming setting. For the sake of simplicity, we assume to be in a regular regression setting and omit the subscript $k$ for the distribution parameter used in the previous subsection. However, we are now interested in the subscript $i=0,1,...,n$, which indicates which data the estimation algorithm has seen already. In the next Section 2.3, where we combine both components, we will employ both subscripts. The issue at hand can be summarized as follows: given some data ${\mathcal{Y}}_{n}$ and ${\mathcal{X}}_{n}$ we have estimated a set of coefficients $\beta_{n}$. We are now interested in updating the coefficients $\beta_{n}$ given a new observation for $Y_{n+1}$ and a new row of $\boldsymbol{X}_{n+1}$. The following paragraphs outline the algorithms for recursive least squares, online LASSO and online LARS. For regularized estimation techniques, the issue of model selection arises. Therefore, the last paragraph discusses online model selection via information criteria.
Recursive Least Squares
We start with the well-known least-squares estimation for the coefficients $\beta^{\text{LS}}$:
$\widehat{\boldsymbol{\beta}}^{\text{LS}}_{n}=({\mathcal{X}}^{\top}_{n}{\mathcal{X}}_{n})^{-1}{\mathcal{X}}^{\top}_{n}{\mathcal{Y}}_{n}.$
given ${\mathcal{Y}}_{n}=(Y_{1},\ldots,Y_{n})^{\top}$ and ${\mathcal{X}}_{n}=(\boldsymbol{X}^{\top}_{1},\ldots,\boldsymbol{X}^{\top}_{n})$. In a repeated batch setting, we can use:
$\widehat{\boldsymbol{\beta}}^{\text{LS}}_{n+1}=({\mathcal{X}}^{\top}_{n+1}{\mathcal{X}}_{n+1})^{-1}{\mathcal{X}}^{\top}_{n+1}{\mathcal{Y}}_{n+1}.$
to calculate $\widehat{\boldsymbol{\beta}}^{\text{LS}}_{n+1}$. Alternatively, employing the Sherman-Morrison formula and calculate $\beta^{\text{LS}}$ via recursive least squares. For an invertible matrix $\boldsymbol{A}$ and vectors $\boldsymbol{u}$ and $\boldsymbol{v}$ it holds
@@
We also have: ${\mathcal{X}}_{n+1}^{\top}{\mathcal{X}}_{n+1}={\mathcal{X}}^{\top}_{n}{\mathcal{X}}_{n}+\boldsymbol{X}_{n+1}^{\top}\boldsymbol{X}_{n+1}$ and therefore
$({\mathcal{X}}_{n+1}^{\top}{\mathcal{X}}_{n+1})^{-1}$ $=({\mathcal{X}}^{\top}_{n}{\mathcal{X}}_{n}+\boldsymbol{X}_{n+1}^{\top}\boldsymbol{X}_{n+1})^{-1}$
$=({\mathcal{X}}^{\top}_{n}{\mathcal{X}}_{n})^{-1}-\frac{({\mathcal{X}}^{\top}_{n}{\mathcal{X}}_{n})^{-1}\boldsymbol{X}^{\top}_{n+1}\boldsymbol{X}_{n+1}({\mathcal{X}}^{\top}_{n}{\mathcal{X}}_{n})^{-1}}{1+\boldsymbol{X}_{n+1}({\mathcal{X}}^{\top}_{n}{\mathcal{X}}_{n})^{-1}\boldsymbol{X}^{\top}_{n+1}}.$ (8)
Thus, we can compute $({\mathcal{X}}_{n+1}^{\top}{\mathcal{X}}_{n+1})^{-1}$ given $({\mathcal{X}}^{\top}_{n}{\mathcal{X}}_{n})^{-1}$ and $\boldsymbol{X}_{n+1}$ without much effort. Furthermore, we receive $\widehat{\boldsymbol{\beta}}^{\text{LS}}_{n+1}$ by noting that ${\mathcal{X}}^{\top}_{n}{\mathcal{X}}_{n}\widehat{\boldsymbol{\beta}}^{\text{LS}}_{n}={\mathcal{X}}^{\top}_{n}{\mathcal{Y}}_{n}$:
$\widehat{\boldsymbol{\beta}}^{\text{LS}}_{n+1}$ $=({\mathcal{X}}^{\top}_{n+1}{\mathcal{X}}_{n+1})^{-1}{\mathcal{X}}^{\top}_{n+1}{\mathcal{Y}}_{n+1}$
$=\widehat{\boldsymbol{\beta}}^{\text{LS}}_{n}+({\mathcal{X}}^{\top}_{n+1}{\mathcal{X}}_{n+1})^{-1}\boldsymbol{X}^{\top}_{n+1}\underbrace{\left(Y_{n+1}-\boldsymbol{X}_{n+1}\widehat{\boldsymbol{\beta}}^{\text{LS}}_{n}\right)}_{\text{forecast error in }n+1}$ (9)
Hence, with (9) and (8), we can efficiently compute expanding window updates for the coefficient vector $\beta^{\text{LS}}$.
Weighted Recursive Least Squares
The RS algorithm presented in the previous Section 2.1 employs a weighted estimation. Hence, we will show in this paragraph how the weighted recursive least squares allow the online estimation of linear models. We define the weight matrix ${\mathcal{W}}$ with known weights for observations $1,...,n,...$ on the diagonal
${\mathcal{W}}_{n}=\begin{pmatrix}w_{1}&0&\dots&0\\ 0&w_{2}&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&w_{n}\end{pmatrix}$ (10)
Under the assumptions that (a) the weights are known and (b) the weights do not change for past observations, the weighted least-squares problem can be written as:
$\widehat{\boldsymbol{\beta}}^{\text{WLS}}_{n}=$ $({\mathcal{X}}^{\top}_{n}{\mathcal{W}}_{n}{\mathcal{X}}_{n})^{-1}{\mathcal{X}}^{\top}_{n}{\mathcal{W}}_{n}{\mathcal{Y}}_{n}$
$=$ $\left(({\mathcal{W}}^{\frac{1}{2}}_{n}{\mathcal{X}}_{n})^{\top}({\mathcal{W}}^{\frac{1}{2}}_{n}{\mathcal{X}}_{n})\right)^{-1}({\mathcal{W}}^{\frac{1}{2}}_{n}{\mathcal{X}}_{n})^{\top}{\mathcal{W}}^{\frac{1}{2}}_{n}{\mathcal{Y}}_{n}$ (11)
by noting that ${\mathcal{W}}_{n}$ is invertible as long as all weights are positive and non-zero. For convenience, we define the vectors
${\mathcal{M}}_{n}$ $=({\mathcal{W}}^{\frac{1}{2}}_{n}{\mathcal{X}}_{n})$ (12)
${\mathcal{M}}_{n+1}$ $=({\mathcal{W}}^{\frac{1}{2}}_{n+1}{\mathcal{X}}_{n+1})$ (13)
$\boldsymbol{M}_{n+1}$ $=\sqrt{w_{n+1}}\boldsymbol{X}_{n+1}$ (14)
and have
${\mathcal{M}}^{\top}_{n+1}{\mathcal{M}}_{n+1}={\mathcal{M}}^{\top}_{n}{\mathcal{M}}_{n}+\boldsymbol{M}^{\top}_{n+1}\boldsymbol{M}_{n+1}$ (15)
so that we can again use the Sherman-Morrison Formula, we can re-write the weighted Gram Matrix $({\mathcal{X}}^{\top}_{n+1}{\mathcal{W}}_{n+1}{\mathcal{X}}_{n+1})^{-1}=({\mathcal{M}}_{n+1}^{\prime}{\mathcal{M}}_{n+1})^{-1}$ for the next observation $n+1$ as
@@
Again, combining (17) and (16) we can efficiently compute the update for the weighted recursive least-squares estimator $\widehat{\boldsymbol{\beta}}^{\text{WLS}}_{n+1}$.
Recursive least squares with exponential forget
The algorithm can be naturally extended to RLS under exponential forgetting
$\hat{\beta}=\arg\min_{\beta}\left\{\left\lVert\gamma^{T-t}(Y-\beta X)\right\rVert^{2}_{2}\right\},$ (18)
where $\gamma$ denotes the forgetting factor. Define the weight matrix $\Gamma_{n}$ as
$\Gamma_{n}=\begin{pmatrix}\gamma^{N-1}&0&\dots&0\\ 0&\gamma^{N-2}&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&\gamma\end{pmatrix}$ (19)
and we note that the exponential discounting of older observations leads to
@@
and the according update equation for the exponentially discounted Gram matrix as
$({\mathcal{X}}^{\top}_{n+1}\Gamma_{n+1}{\mathcal{X}}_{n+1})^{-1}=\frac{1}{\gamma}\left(({\mathcal{X}}^{\top}_{n}\Gamma_{n}{\mathcal{X}}_{n})^{-1}-\frac{({\mathcal{X}}^{\top}_{n}\Gamma_{n}{\mathcal{X}}_{n})^{-1}\boldsymbol{X}^{\top}_{n+1}\boldsymbol{X}_{n+1}({\mathcal{X}}^{\top}_{n}\Gamma_{n}{\mathcal{X}}_{n})^{-1}}{\gamma+\boldsymbol{X}_{n+1}({\mathcal{X}}^{\top}_{n}\Gamma_{n}{\mathcal{X}}_{n})^{-1}\boldsymbol{X}^{\top}_{n+1}}\right).$ (21)
Similar to the previous paragraph, we can efficiently update the coefficients by combining Equations (17) and (21).
@@
The LASSO estimator for linear models minimizes:
$\widehat{\beta}_{n}=\arg\min_{\beta}\left\{\left\|{\mathcal{W}}^{\frac{1}{2}}\left({\mathcal{Y}}_{n}-\beta{\mathcal{X}}_{n}\right)\right\|_{2}^{2}+\lambda_{n}\left\|\beta\right\|_{2}\right\}.$
@@
Friedman et al. (2007, 2010) have introduced Path-wise cyclic coordinate descent (CCD) for LASSO, ridge and elastic net regression problems, and it remains the computationally efficient method to estimate coefficients. For a given regularization parameter $\lambda$, we repeatedly update the coefficient vector $\boldsymbol{\beta}$ of length $J$ by taking
$\beta_{j}\xleftarrow{}\frac{S\left(\sum_{n=1}^{N}w_{n}x_{n,j}(y_{n}-\tilde{y}_{n}^{(j)}),\lambda\right)}{\sum_{n=1}^{N}w_{n}x_{n,j}^{2}}$ (22)
@@
where $S(\hat{\beta},\lambda)=\text{sign}(\hat{\beta})(\lvert\hat{\beta}\rvert-\lambda)_{+}$ is the soft-thresholding function. Note that, implicitly, Friedman et al. (2010) assume that $\sum_{n=1}^{N}w_{n}=1$. We use a numerical convergence criterion to break the algorithm after convergence. As proposed in Angelosante et al. (2010) and Messner & Pinson (2019), we rewrite the update Equation 22 to calculate the updated parameter from the Gramian matrix ${\mathcal{G}}={\mathcal{X}}^{\top}{\mathcal{X}}$ and the $y$-Gram matrix ${\mathcal{H}}={\mathcal{X}}^{\top}{\mathcal{Y}}$. In the univariate case, we can write the online coordinate descent (OCD) update rule for the LASSO estimator as:
@@
$\beta_{j}\xleftarrow{}\frac{S\left({\mathcal{H}}\left[j\right]-{\mathcal{G}}\left[j,:\right]\boldsymbol{\beta}+{\mathcal{G}}\left[j,j\right]\beta_{j},\lambda\right)}{{\mathcal{G}}\left[j,j\right]}$ (23)
@@
Therefore, the algorithm only needs to store the $J\times J$ matrix ${\mathcal{G}}$ and the $J\times 1$ vector ${\mathcal{H}}={\mathcal{X}}^{\top}{\mathcal{Y}}$. Similar to the classical coordinate descent, we can run the algorithm for a decreasing sequence of regularization strengths $\lambda$, starting with $\lambda_{\text{max}}=\max{}\lvert{\mathcal{X}}^{\top}{\mathcal{Y}}\rvert$ as the element-wise maximum in ${\mathcal{H}}$ and using an exponential grid towards $\lambda_{\text{min}}=\epsilon\lambda_{\text{max}}$ with $\epsilon_{\lambda}=0.001$ as typical values (Friedman et al., 2010, 2007; Messner & Pinson, 2019). Algorithm 1 presents the algorithm schematically.
Data: $\boldsymbol{X}_{n+1},Y_{n+1},w_{n+1}$
Result: $\boldsymbol{\beta}_{n+1}$
Update the Gramian matrices ${\mathcal{G}}_{n+1}$ and ${\mathcal{H}}_{n+1}$ according to Equation 16.
for $\lambda$ do
       Set starting coefficients $\boldsymbol{\beta}_{\lambda}\leftarrow\boldsymbol{\beta}_{\lambda[-1]}$
       forall $j\in 1,...,J$ do
             update $\beta_{j,\lambda}$ according to Equation 23

### v2

@@ 2.1 Batch GAMLSS
8                   Regress $\boldsymbol{z}_{k}^{[m,r]}$ against ${\mathcal{X}}_{k}$ using weights $\boldsymbol{W}_{k,k}^{[m,r]}$ to obtain $\hat{\boldsymbol{\beta}}_{k}^{[m,r+1]}$.
9                   Calculate the updated $\hat{\boldsymbol{\eta}}_{k}^{[m,r+1]}$ and $\hat{\boldsymbol{\theta}}_{k}^{[m,r+1]}$.
10                   Evaluate the convergence.
11
12            End the inner cycle on the convergence of $\hat{\boldsymbol{\beta}}_{k}^{[m,...]}$
13             Set $\hat{\boldsymbol{\beta}}_{k}^{[m+1,0]}\leftarrow\hat{\boldsymbol{\beta}}_{k}^{[m,...]}$, set $\hat{\boldsymbol{\eta}}_{k}^{[m+1,0]}\leftarrow\hat{\boldsymbol{\eta}}_{k}^{[m,...]}$, set $\hat{\boldsymbol{\theta}}^{[m+1,0]}\leftarrow\hat{\boldsymbol{\theta}}_{k}^{[m,...]}$.
14
15      End the outer cycle if the change in the penalized likelihood is sufficiently small.
16
Output: $\widehat{\boldsymbol{\beta}}_{k,n}$ and $\widehat{\boldsymbol{\theta}}_{k,n}$
Algorithm 1 Batch GAMLSS (RS-Algorithm), see Rigby and Stasinopoulos (2005)
Relationship between GAMLSS and IRLS
It is fairly clear that a GAMLSS approach which models only the location and scale parameter of a specific parametric distribution is to some extent related to a mean-variance modeling approach by iteratively reweighted least squares algorithm (IRLS) for conditional heteroskedasticity (Dette and Wagener, 2013; Ziel et al., 2016; Ziel, 2016; Francq and Zakoian, 2019). In the latter, the variance equation is estimated by regressing the input variables on the squared residuals $({\mathcal{Y}}-\widehat{\mu})^{2}$, where $\widehat{\mu}$ are the predictions of the mean model. This general approach is commonly used in ARMA-GARCH-type modelling frameworks where the mean equation residuals ${\mathcal{Y}}-\widehat{\mu}$ serve as a proxy for the variance relationship (Francq and Zakoian, 2019, Chapter 6). However, the relationship between GAMLSS and IRLS approaches for mean-variance models is rather complicated. It depends substantially on the parametric distribution ${\mathcal{D}}$, including its parameterization for the location and scale parameters $\theta_{1}$ and $\theta_{2}$ and on the link function $g_{1}$ and $g_{2}$. Nevertheless, if we consider the normal distribution in a GAMLSS framework and parameterize $\mu$ and $\sigma^{2}$ with identity link function in Equation 2, we receive a setting that is closely aligned with the mean-variance IRLS approach:
Proposition 1
When considering a GAMLSS model for the normal distribution with density $f(x|\mu,\sigma)=1/\sqrt{2\pi\sigma^{2}}\exp\left(-(x-\mu)^{2}/\sigma^{2}\right),$ a parameterization of $\theta_{1}=\mu$ and $\theta_{2}=\sigma^{2}$ and identity link functions $g_{1}(\theta)=\theta$ and $g_{2}(\theta)=\theta$ then the RS algorithm with the Newton-Raphson scoring has the same working vectors as the iteratively reweighted least squares algorithm (IRLS) for mean-variance in Dette and Wagener (2013); Ziel (2016), i.e. it holds
$z_{1}=z_{\mu}=$ $y,$ $z_{2}=z_{\sigma^{2}}$ $=(y-\mu)^{2}.$ (8)
Moreover, the GAMLSS approach is updated using the weights
$\boldsymbol{W}_{1,1}=$ $\dfrac{1}{\sigma^{2}},$ $\boldsymbol{W}_{2,2}=$ $\dfrac{1}{2\sigma^{4}}.$ (9)
Hence, when choosing the constant weights in the mean equation equal to $1/\sigma^{2}$ and in the variance equation equal to $1/(2\sigma^{4})$ in the IRLS algorithm, then it is equivalent to the GAMLSS approach in the setting of the proposition. However, Ziel (2016) proposed the IRLS algorithm with $1/\sigma^{\delta}$ for the mean equation (corresponding to the GAMLSS algorithm above for $\delta=2$) and constant weights in the variance equation.
Error Propagation
The working vectors given in Equation 8 in Proposition 1 also underscore an important fact in distributional modelling: The model for the conditional mean $\hat{\mu}$ is part of the model for the conditional scale $\hat{\sigma}$. Therefore, a low-quality model for $\hat{\mu}$ will propagate through the estimation and lead to an overestimation of the conditional variance (Ziel, 2022) and hard-to-predict effects on higher moments. This indicates that higher moment parameters tend to require more care when dealing with overfitting, e.g. choosing a more conservative regularization framework.
@@ 2.2 Online estimation of regularized linear models
Let us now move to the online or streaming setting. For the sake of simplicity, we assume to be in a regular regression setting and omit the subscript $k$ for the distribution parameter used in the previous subsection. However, we are now interested in the subscript $i=0,1,...,n$, which indicates which data the estimation algorithm has seen already. In the next Section 3, where we combine both components, we will employ both subscripts. The issue at hand can be summarized as follows: given some data ${\mathcal{Y}}_{n}$, ${\mathcal{X}}_{n}$ and weights ${\mathcal{W}}_{m}=\operatorname{diag}\left(w_{1},...,w_{n}\right)$ we have estimated a set of coefficients $\hat{\beta}_{n}$. We are now interested in updating the coefficients $\hat{\beta}_{n}$ given a new observation for $Y_{n+1}$, a new row of $\boldsymbol{X}_{n+1}$ and a new weight $w_{n+1}$. The following paragraphs outline the algorithms for (exponentially discounted) weighted recursive least squares and online LASSO. For regularized estimation techniques, the issue of model selection arises. Therefore, the last paragraph discusses online model selection via information criteria.
Weighted Recursive Least Squares
The following introduction is a condensed textbook introduction (e.g. Haykin, 2014, Chapter 10-12). Nevertheless, it paves the ground for the online LASSO and we emphasize a few interesting points for the online distributional learning algorithm introduced in Section 3. The weighted least squares regression problem estimates the coefficients $\hat{\beta}^{\text{WLS}}$ that minimize the loss:
$\hat{\beta}^{\text{WLS}}=\arg\min_{\beta}\left\{\left\lVert{\mathcal{W}}^{\frac{1}{2}}_{n}({\mathcal{Y}}_{n}-\beta{\mathcal{X}}_{n})\right\rVert^{2}_{2}\right\},$ (10)
by noting that ${\mathcal{W}}_{n}$ is invertible as long as all weights are positive and non-zero, we can write the classic OLS estimator
$\widehat{\boldsymbol{\beta}}^{\text{WLS}}_{n}=({\mathcal{X}}^{\top}_{n}{\mathcal{W}}_{n}{\mathcal{X}}_{n})^{-1}{\mathcal{X}}^{\top}_{n}{\mathcal{W}}_{n}{\mathcal{Y}}_{n}=\left(({\mathcal{W}}^{\frac{1}{2}}_{n}{\mathcal{X}}_{n})^{\top}({\mathcal{W}}^{\frac{1}{2}}_{n}{\mathcal{X}}_{n})\right)^{-1}({\mathcal{W}}^{\frac{1}{2}}_{n}{\mathcal{X}}_{n})^{\top}{\mathcal{W}}^{\frac{1}{2}}_{n}{\mathcal{Y}}_{n}$
Under the assumptions that (1) the weights are known and (2) do not change for past observations, the weighted least-squares problem can be re-formulated as a recursive algorithm for the next observation $n+1$. For convenience, we define the vectors ${\mathcal{M}}_{n}={\mathcal{W}}^{\frac{1}{2}}_{n}{\mathcal{X}}_{n}$, ${\mathcal{M}}_{n+1}={\mathcal{W}}^{\frac{1}{2}}_{n+1}{\mathcal{X}}_{n+1}$, and $\boldsymbol{M}_{n+1}=\sqrt{w_{n+1}}\boldsymbol{X}_{n+1}$. The Sherman-Morrison Formula states that for an invertible matrix $\boldsymbol{A}$ and vectors $\boldsymbol{u}$ and $\boldsymbol{v}$ it holds
@@
Since we have ${\mathcal{M}}^{\top}_{n+1}{\mathcal{M}}_{n+1}={\mathcal{M}}^{\top}_{n}{\mathcal{M}}_{n}+\boldsymbol{M}^{\top}_{n+1}\boldsymbol{M}_{n+1}$ we can rewrite the inverted weighted Gram Matrix $({\mathcal{X}}^{\top}_{n+1}{\mathcal{W}}_{n+1}{\mathcal{X}}_{n+1})^{-1}=({\mathcal{M}}_{n+1}^{\prime}{\mathcal{M}}_{n+1})^{-1}$ for the next observation $n+1$ as
@@
By combining (12) and (11) we can efficiently compute the update for the weighted recursive least-squares estimator $\widehat{\boldsymbol{\beta}}^{\text{WLS}}_{n+1}$. An important special case is the recursive least squares with exponential forget, which minimizes the loss
$\hat{\beta}=\arg\min_{\beta}\left\{\left\lVert(\gamma^{N-n})^{\frac{1}{2}}({\mathcal{Y}}_{n}-\beta{\mathcal{X}}_{n})\right\rVert^{2}_{2}\right\},$ (13)
where $\gamma=1-\varphi$ and $\varphi$ denotes the forgetting factor. We define the weight matrix $\Gamma_{n}=\operatorname{diag}\left(\gamma^{N-1},\gamma^{N-2},...,\gamma^{1},\gamma^{0}\right)$ and note that the exponential discounting of older observations leads to
@@
and the according update equation for the exponentially discounted inverted Gram matrix
$({\mathcal{X}}^{\top}_{n+1}\Gamma_{n+1}{\mathcal{X}}_{n+1})^{-1}=\frac{1}{\gamma}\left(({\mathcal{X}}^{\top}_{n}\Gamma_{n}{\mathcal{X}}_{n})^{-1}-\frac{({\mathcal{X}}^{\top}_{n}\Gamma_{n}{\mathcal{X}}_{n})^{-1}\boldsymbol{X}^{\top}_{n+1}\boldsymbol{X}_{n+1}({\mathcal{X}}^{\top}_{n}\Gamma_{n}{\mathcal{X}}_{n})^{-1}}{\gamma+\boldsymbol{X}_{n+1}({\mathcal{X}}^{\top}_{n}\Gamma_{n}{\mathcal{X}}_{n})^{-1}\boldsymbol{X}^{\top}_{n+1}}\right)$ (15)
which allows for an efficient online update of the OLS coefficients under exponential forgetting. The weighted least squares allows to combine user-defined sample weights and exponential forget. For the exponentially discounted, weighted Gram matrix ${\mathcal{G}}_{n}=({\mathcal{X}}_{n}^{\top}\Gamma_{N}{\mathcal{W}}_{n}{\mathcal{X}}_{n})$, we have
${\mathcal{G}}_{n+1}=\gamma{\mathcal{G}}_{n}+w_{n+1}\boldsymbol{X}^{\top}_{n+1}\boldsymbol{X}_{n+1}$ (16)
and accordingly the update of the inverted Gram matrix ${\mathcal{G}}_{n+1}^{-1}$ can be written as:
${\mathcal{G}}^{-1}_{n+1}=\dfrac{1}{\gamma}\left({\mathcal{G}}^{-1}_{n}-\frac{w_{n+1}{\mathcal{G}}^{-1}_{n}\boldsymbol{X}^{\top}_{n+1}\boldsymbol{X}_{n+1}{\mathcal{G}}^{-1}_{n}}{\gamma+w_{n+1}\boldsymbol{X}_{n+1}{\mathcal{G}}^{-1}_{n}\boldsymbol{X}^{\top}_{n+1}}\right)$ (17)
which can be plugged into Equation 11 to update the coefficients $\hat{\beta}$, as in Algorithm 2.
Input: New observations $\boldsymbol{X}_{n+1},Y_{n+1},w_{n+1}$ and stored ${\mathcal{G}}^{-1}_{n},\hat{\beta}_{n}$.
1 Update ${\mathcal{G}}^{-1}_{n}\rightarrow{\mathcal{G}}^{-1}_{n+1}$ according to Equation 11 resp. 17.
2 Update $\hat{\beta}_{n}+{\mathcal{G}}^{-1}_{n+1}w_{n+1}\boldsymbol{X}^{\top}_{n+1}\left(Y_{n+1}-\boldsymbol{X}_{n+1}\hat{\beta}_{n}\right)\rightarrow\hat{\beta}_{n+1}$ (see Equation 12)
Output: $\hat{\beta}_{n+1}$ and store updated Gramian ${\mathcal{G}}^{-1}_{n+1}$.
Algorithm 2 Recursive Weighted Exponentially Discounted Least-Squares
@@
The well-known LASSO estimator for linear models minimizes a weighted combination of the sum-of-squared residuals and a $\ell_{1}$-penalty term on the coefficient vector
$\hat{\beta}_{n}=\arg\min_{\beta}\left\{\left\|{\mathcal{W}}^{\frac{1}{2}}\left({\mathcal{Y}}_{n}-\beta{\mathcal{X}}_{n}\right)\right\|_{2}^{2}+\lambda_{n}\left\|\beta\right\|_{1}\right\}.$
@@
to achieve sparsity in the estimated coefficients. Friedman et al. (2007, 2010) have introduced path-wise cyclic coordinate descent (CCD) for LASSO, ridge and elastic net regression problems, and it remains the computationally efficient method to estimate coefficients. For a given regularization parameter $\lambda$, we repeatedly update the coefficient vector $\boldsymbol{\beta}$ of length $J$ by taking
$\hat{\beta}_{j}\xleftarrow{}\frac{S\left(\sum_{n=1}^{N}w_{n}x_{n,j}(y_{n}-\tilde{y}_{n}^{(j)}),\lambda\right)}{\sum_{n=1}^{N}w_{n}x_{n,j}^{2}}$ (18)
@@
where $S(\hat{\beta},\lambda)=\text{sign}(\hat{\beta})(\lvert\hat{\beta}\rvert-\lambda)_{+}$ is the soft-thresholding function. Note that, implicitly, Friedman et al. (2010) assume that $\sum_{n=1}^{N}w_{n}=1$. We use a numerical convergence criterion to break the algorithm after convergence. As proposed in Angelosante et al. (2010) and Messner and Pinson (2019), we rewrite the update Equation 18 to calculate the updated parameter from the Gramian matrix ${\mathcal{G}}={\mathcal{X}}^{\top}{\mathcal{X}}$ and the $y$-Gram matrix ${\mathcal{H}}={\mathcal{X}}^{\top}{\mathcal{Y}}$ and their potentially exponentially discounted and/or weighted counterparts (see Section 2.2 and Equation 16). Hence, we can write the online coordinate descent (OCD) update rule for the LASSO estimator as
@@
$\hat{\beta}_{j}\xleftarrow{}\frac{S\left({\mathcal{H}}\left[j\right]-{\mathcal{G}}\left[j,:\right]\boldsymbol{\beta}+{\mathcal{G}}\left[j,j\right]\hat{\beta}_{j},\lambda\right)}{{\mathcal{G}}\left[j,j\right]}.$ (19)
@@
Therefore, the algorithm only needs to store the $J\times J$ matrix ${\mathcal{G}}$ and the $J\times 1$ vector ${\mathcal{H}}$. Similar to the classical coordinate descent, we can run the algorithm for a decreasing sequence of regularization strengths $\lambda$, starting with $\lambda_{\text{max}}=\max{}\lvert{\mathcal{H}}\rvert$ as the element-wise maximum in ${\mathcal{H}}$ and using an exponential grid towards $\lambda_{\text{min}}=\epsilon\lambda_{\text{max}}$ with $\epsilon_{\lambda}=0.001$ as typical values (Friedman et al., 2010, 2007; Messner and Pinson, 2019). Algorithm 3 presents the algorithm schematically.
Input: New observations $\boldsymbol{X}_{n+1},Y_{n+1},w_{n+1}$ and stored ${\mathcal{G}}_{n},{\mathcal{H}}_{n}$.
1 Update ${\mathcal{G}}_{n}\rightarrow{\mathcal{G}}_{n+1}$ and ${\mathcal{H}}_{n}\rightarrow{\mathcal{H}}_{n+1}$ according to Equation 16.
2 Update $\lambda_{\text{max}}=\max{}\lvert{\mathcal{G}}_{n+1}\rvert$ and initialize $\boldsymbol{\lambda}$ as exponential grid.
3 for $\lambda\in\boldsymbol{\lambda}$ do
4       Set starting coefficients $\boldsymbol{\beta}_{\lambda}\leftarrow\boldsymbol{\beta}_{\lambda[-1]}$
5       while not converged do
6             forall $j\in 1,...,J$ do
7                   Update $\hat{\beta}_{j,\lambda}$ according to Equation 19
8